By 15:00 Friday the LHC fully closed and ready for beam . At this time access was required in the PS for a faulty septum. Beam interlock tests finished.
At 17:30 the single bunch low emittance probe beam was taken ion to the TI2 TED (dump block) at the end of the TI2 transfer line. In the LHC the septa magnet were pulsing and the cold magnets cycled. The LHC injection kickers were timed in with the beam
18:54 Kickers off, TED taken out, after a couple of corrections at the end of TI2 beam was seen in the LHC for the first time. At this point the dump block [TDI] just after the injection point was in place. Beam loss monitors, Beam position monitors, vacuum guages and screens as witness. Much excitement in control room.
Some time taken to syncchronize the RF and timing control systems.
21:42 TDI out - beam through Alice and to the end of sector 23 first shot. Beam on screen. Trajectory shortly after.
23:30 Trajectory to IR3 corrected with a peak in the arc less than 3 mm.
Kick response measurements revealed phase error in arc. All corrector polarites correct. One BPM with wrong polarity.
02:35 Quench Attempt to probe the aperture. Quench on MB.A8.L3 with an oscillation amplitude of ~ 12 mm from MCBV.9R2.B1 (80 murad). At the same time a huge number of corrector circuits go off. QPS trigger time 2:19:48:597 and last BPM acquisition is 2:19:48. The trim on the corrector was entered at 2:19:11.
Quench proved to be a gentle one but sets quench limit for prompt loss to be around 4 10^9 protons - in almost exact agreement with earlier predictions:
"The intensity of the bunch shall therefore not be much larger than 3 10^9 protons."
Magnet temperature profile during quench
Quench as seen by beam loss monitors
LHC put to bed until the morning.
Access required Saturday to fixed a quench heater power supply.
Beam back to IR3 after re-cycling at around 15:00 - straight to IR3 - things look very reproducible. Some spent checking RF anfd beam instrumentation timing before starting aperture measurements in the injection region.
- the aperture at MSI looks just as expected - we start seeing losses at about +/-6 nominal sigma in vertical plane, no losses seen up to 7 nominal sigma in h plane.
No chance yet to scan further - not really a priority - the aperture at Q5 for injected beam looks very tight (again as expected) - predicted by JBJ and JU to be about 5 sigma towatrds the bottom at the entrance of the Q5 (which has been tilted to even get this much)
- and, yep, we started seeing significant losses at this setting. The losses were definately on the Q5 (BLMs lit up and also confirmed from the injection region screen shots).
- some results still to understand, where it appears we might just be scraping on Q4 with the MKI off - this was also a concern - we had no time to look more in detail at this. - Q5 doesn't quench with 5e9.
- MKI's don't like beam loss much - we need to make sure we do a SoftStart whenever there is beam loss in this region - will have to see with JW and VK how to enforce this.
Scattered particles seemed to have caused a flash-over in the injection kickers - beam off while kickers recovered.
22:00 beam back to IR2. 22:30 collimators left of IP3 out.
23:05 Beam through IR3 - things look good. BPMs again give data first time.
23:00 to 07:00 Beautiful set of aperture measurements performed. All trajectories from vertical aperture measurements on one plot.
09:00 Commissioning of RF frequency SPS/LHC leading into dispersion measurements in LHC. First measurement around 13:55.
LHC BPM system intensity threshold measured to be ~1.5e+9 by varying the intensity coming from the PS (extr. intensity from 1e+9). (c/o Rhodri Jones)
Matching screen beam images.
At 19:30 TDI in for finish the aperture check of the injection region.
22:30 Injection aperture checks finished. Re-cycle.
23:15 Loss of power power permit on a lots of circuit...
00:40 LHC back - apparently main quads heaters fired for some reason
Wrestling with LHC injeciton kickers... it seems like regular operations with Bernd & sons waiting patiently to do some measurements.
|At the end of the aperture measurements in the injection region, the kicker was re-conditioned with the SlowStart. When it was ready, a problem of the QPS system caused the lost of powering conditions for RQD circuit, which triggered the loss of all the arc circuits. See detailed description of what happened in the corresponding entry above.
By the time the magnets were back in operation, we needed to re-condition the kicker again. After that we could not start the beam operation because we could not actually pulse the kicker. After some struggling we figured out that the reason for this problem was that we did not propely requeste the pulsing 'Enable' from the kicker control software.
Beam was back and we had about a couple of hours of beam time that we used to perform the following measurements (details of which are outlined above):
- BLM studies in the collimator region
- Orbit response studies with different polarities of the RQTL11.L3B1
LHC off at 06:00
|Injection quality checks|
|Timing of BST & pre-pulses|
|Reduced intensity from PS|
|BLMs - in depth|
|Warm PM calibration|